Two rainstorms occurred in one week in a certain area. In the first rainstormn 15ML of rain fell per hour, and in the second rainstorm 40ML of rain fell per hour. Rain fell that week for a total of 65 hours for a total rainfall of 1850ML. How long was each of the two rainstorms?
Answer by stanbon(75887)

Two rainstorms occurred in one week in a certain area. In the first rainstormn 15ML of rain fell per hour, and in the second rainstorm 40ML of rain fell per hour. Rain fell that week for a total of 65 hours for a total rainfall of 1850ML. How long was each of the two rainstorms?
------------------------
Let the length of the 40ML/hr rain be "x" hr. Amt of rain is 40x ML
---------------------
The length of the 15ML/hr rain is 65-x hrs; rain amt= 15(65-x)= 975-15x ML
----------------------
rain + rain = 1850 ML
40x + 975-15x = 1850
25x = 875
x= 35 hr (length of the 40ML/hr rain)
65-15=30 hr (length of the 15ML/hr rain)
====================
Cheers,
Stan H.